OpenStudy (anonymous):

I need a refresher on sig figs (Significant Digits). Short explanation and maybe a few examples would be amazing.

OpenStudy (anonymous):

When multiplying two numbers the smaller amount of significant digits is used. This applies for division also. \[1.5*1.25=1.9\] \[\frac{1.5}{1.125}=1.3\] When adding and subtracting the least number of decimal places is what matters for example: \[1.1-1.005=0.1\] The key thing to recall is leading zeros do not matter when considering SFs, but trailing zeros are significant. So 0.001 has 1 SF where as 0.0010 has 2 SF. This is despite the numbers having the same value.

OpenStudy (anonymous):

1. significant figures are always count from non zero digit for exampl-0.00123 here SF=3 1.000003 sf=7 2.addition of significant figure 1.003 +0.12 _______ 1.123 wen we do addition we take only dat much numbers see in the example here 0.12 has sf=2 and 1.003has sf=4 so in the result we write the ans acc. to least SF. so in the result 1.12:3 we will aprox {rule-wen the nxt digit is above from 5 den we will add one digit before of it and if the digit is less than 5 we will not change the num which is before it} now here in the ans 3 is less than the 5 then the aprox will be = 1.22 and SF is3 one more example 2.456 +1.45 _______ 1.9 06 now least significant digit is 3 so we will do approx till 3 digits means till 0 here 6 is larger dan 5 so we will add 1 its before digit dat is 0 and it will become 1 so the result will be =1.91 ans SF=3 3.multiplication it is v.easy u have to just calculate the least SF and the ans will be automatically have same SF digit example=1.23*2.1=? they never ask the calculation dey always ask the SF in ans so here leat SF =2 therefore in result it will have SF=2 4.division it is also easy u have to jst find the least SF in the ques example= 1.34-4.5 _____________ 123.5 *10^5 here 4.5 has least SF =2 so in the ans total SF will be 2 NOTE=1. IN CALCULATION WE NEVER COUNT WHOLE NUMBER (2,3,ETC) WE ONLY TAKE DECIMAL POINT NUMBERS 2.IF THE NUMBER IS IN POWER DAT NUMBER IS ALSO NT TAKEN IN CALCULATION JST IGNORE THEM.. :-)